Patel and Sons Inc. uses a standard cost system to apply factory overhead costs to units produced. Practical capacity for the plant is defined as 50,100 machine hours per year, which represents 25,050 units of output. Annual budgeted fixed factory overhead costs are $250,500 and the budgeted variable factory overhead cost rate is $2.00 per unit. Factory overhead costs are applied on the basis of standard machine hours allowed for units produced. Budgeted and actual output for the year was 18,500 units, which took 39,100 machine hours. Actual fixed factory overhead costs for the year amounted to $246,100 while the actual variable overhead cost per unit was $1.90.

Based on the information provided above, calculate the following factory overhead variances for the year. Indicate whether each variance is favorable (F) or unfavorable (U). (Do not round intermediate calculations. Round your final answers to nearest whole dollar amount.)

(a) Total overhead variance Unfavorable
(b) Total flexible-budget variance Favorable
(c) Production volume variance $65,500 Unfavorable
Based on the information provided above, provide an appropriate end-of-year closing entry for each of the following two independent situations: (a) the net factory overhead cost variance is closed entirely to Cost of Goods Sold (CSG), and (b) the net factory overhead variance is allocated among WIP Inventory, Finished Goods Inventory, and CGS using the following percentages: 20%, 20%, and 60%, respectively. (Do not round intermediate calculations. Round your final answers to nearest whole dollar amount. If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)
Answer #1
Answer
Standard time for producing one unit of output = 50100 / 25050 = 2 hours per unit of output
Budgeted Fixed Overhead = $250500  
Budgeted overhead absorption rate per hour = $2,50,500 / 50100 hours = $5 per hour.
Budgeted units produced during the year was 18500 units. Standard hours for 18500 units = 2 x 18500 = 37000 hours.
Answer A Total Fixed Overhead Variance = standard hours allowed for actual production x standard overhead absorption rate per hour - Actual Fixed overhead cost incurred
37000 hrs. x $5 per hour - $2,46100 = $61,100 (U).
Standard Variable Overhead rate = $2 per unit.
Variable Overhead Variance = Actual output x Standard rate per unit - Actual variable overhead
18500 units x $ 2. per unit – 18500 units x $1.9 per unit = 1,850 (F).
Total Overhead Variance = Total Fixed Overhead Variance + Total Variable Overhead Variance = $61100 (U) + $1850 (F) = $59250(U).
Answer B When flexible budget will be prepared for actual production of 18500 units, the fixed expense will remain same as the budget i.e. $2,61,000.
Flexible overhead budget total for actual output of 18500 units = $2,50500 + 18500 x $2
$213500
Actual overhead expense = $246100+ 18500 x $1.9 = $210950
Total Flexible budget variance = $213500-$ 210950 = $2550(F).
Answer C Production Volume Variance
(Budgeted Hours - Standard Hours required for actual production) x Standard overhead rate per hour = (50100 hrs – 18500 units x 2 hrs /unit) x $5/hr = $65500
Please hit LIKE button if this helped.